OpenStudy (anonymous):

15squareroot6 over 3squareroot5

OpenStudy (anonymous):

\[15\sqrt{6}\over3\sqrt{5}\]

OpenStudy (anonymous):

\[\frac{ 15\sqrt{6} }{3\sqrt{5} }*\frac{ \sqrt{5} }{\sqrt{5} }=\frac{ 15\sqrt{30} }{ 3*5 }=\sqrt{30}\]

OpenStudy (anonymous):

\[\sqrt{5}*\sqrt{5}=\left( \sqrt{5} \right)^{2}=5\]

OpenStudy (anonymous):

\[\frac{ 15\sqrt{6} }{3\sqrt{5} }*\frac{ \sqrt{5} }{\sqrt{5} }=\frac{ 15\sqrt{30} }{ 3*5 }=\sqrt{30}\]